The revenue for a sandwich shop is directly proportional
to its advertising budget. When the owner spent $2000 a month on advertising, the revenue was $120,000. If the revenue is now $180,000, how much is the owner
spending on advertising?

Answer by stanbon(48552)

The revenue for a sandwich shop is directly proportional
to its advertising budget. When the owner spent $2000 a month on advertising, the revenue was $120,000. If the revenue is now $180,000, how much is the owner
spending on advertising?
---------------
Let "x" be the unknown advertising amt.
EQUATION:
x/180,000 = 2000/120000
x = 180000(1/60)
x =$3,000
===========
Cheers,
Stan H.